### v1

# Generative AI-Powered Plugin for Robust

Abstract
Federated learning enables edge devices to collaboratively train a global model while maintaining data privacy by keeping data localized. However, the Non-IID nature of data distribution across devices often hinders model convergence and reduces performance. In this paper, we propose a novel plugin for federated optimization techniques that approximates Non-IID data distributions to IID through generative AI-enhanced data augmentation and balanced sampling strategy. Key idea is to synthesize additional data for underrepresented classes on each edge device, leveraging generative AI to create a more balanced dataset across the FL network. Additionally, a balanced sampling approach at the central server selectively includes only the most IID-like devices, accelerating convergence while maximizing the global model’s performance. Experimental results validate that our approach significantly improves convergence speed and robustness against data imbalance, establishing a flexible, privacy-preserving FL plugin that is applicable even in data-scarce environments.
Index Terms: distributed learning, federated learning, privacy-preserved
I Introduction
Deep learning [1] has shown remarkable performance in various domains by learning complex representations from large datasets. However, achieving such performance often requires extensive data, which is impractical in many real-world scenarios due to privacy concerns and the prohibitive costs associated with data sharing [2]. Federated learning (FL) [3] addresses these challenges by enabling edge devices to train models locally and share only the updated models with a central server, thus maintaining data privacy and security [4]. FL, therefore, provides a crucial solution for privacy-preserving learning, especially in sensitive sectors such as healthcare.
While FL achieves comparable performance to centralized learning by aggregating models from local devices, it often requires numerous global epochs between the central server and the edge devices to achieve satisfactory accuracy [5]. This challenge intensifies when data at each edge device is non-independent and identically distributed (Non-IID) [6] or when computational capabilities vary across devices [7]. Such statistical and computational heterogeneity among edge devices often results in local models drifting towards distinct minima, thus degrading the performance of the aggregated model [8] and increasing the communication overhead required for convergence [9], [10]. Addressing these issues is essential for deploying FL effectively in real-world applications, particularly in medical tasks where data distributions are often highly imbalanced [11].
Figure 1: Illustration of FL in a heterogeneous environment with the proposed plugin: This approach approximates Non-IID data distributions to IID, balancing statistical heterogeneity across edge devices to enhance convergence.
Several methods have been proposed to address this client-drift phenomenon by modifying the local training process. For instance, proximal regularization [12] and restricted-softmax function [13] help alleviate client-drift but increase edge-side complexity. In this paper, we propose a novel plugin to mitigate client drift in heterogeneous FL environments. Our approach involves paraphrasing data using generative AI and selecting an optimal set of edge devices to participate in the updates, thereby approximating the Non-IID data distributions to IID as shown in Fig. 1. Our proposed plugin provides a flexible solution that can be adapted to a wide range of FL frameworks and practical use cases, especially relevant in medical text classification.
Key idea is to utilize generative AI to synthesize data, balancing the distribution of data across edge devices to approximate IID conditions. By addressing data imbalance and utilizing balanced sampling at the central server, we reduce the need for prolonged communication rounds and improve convergence speed. To the best of the authors’ knowledge, this is the first approach to apply generative AI for balancing Non-IID distributions within FL, specifically targeting heterogeneous environments with computational and data distribution challenges. Our method thus represents a significant advance in overcoming core limitations in FL. Main contributions of this paper are as follows:
  • We propose a generative AI-based method for alleviating Non-IID data distributions, referred to as Non-IID to IID Approximation, applicable to various FL schemes.
  • We use generative AI at the edge device level to synthesize data, filling gaps in underrepresented classes to achieve a more balanced distribution without additional real data.
  • We design a balanced sampling approach at the central server to replace random sampling, selecting the most statistically balanced edge device set for aggregation to reduce communication rounds and accelerate model convergence.
  • We demonstrate the effectiveness of this approach by applying it to representative FL methods showing improved training time and accuracy on medical text classification tasks under heterogeneous setting.
The remainder of this paper is organized as follows. In section II, we briefly introduce representative FL methods. In section III, we describe the problem and system model of the proposed method. In section IV, the performance of the FL techniques are validated through numerical results. Finally, we present our concluding remarks in section V.
II Preliminaries
In this section, we summarize the commonly used FL methods used in distributed training, namely FedAvg, FedProx, and FedRS. These methods address challenges in model aggregation and client variability in FL. We denote $N$ as the number of participating edge devices and $g_{e}$ as the global epoch for aggregation.
FedAvg: FedAvg is the foundational FL method where each edge device independently trains a model on its local data and then sends the updated parameters to the central server. Then, the central server averages these locally updated models to create a new global model, which is then redistributed to the edge devices for further training. FedAvg approach is particularly effective in environments where data distributions across devices are homogeneous. However, FedAvg may experience challenges, such as client drift, in Non-IID settings where data distributions vary significantly among edge devices, leading to slower convergence and potential instability.
FedProx: FedProx extends FedAvg by introducing a proximal term at edge device local learning process that helps alleviate client-drift in heterogeneous environments. By constraining each device’s update to be close to the current global model, the proximal term reduces the impact of divergent local models. Specifically, each edge device $n$ minimizes the following objective during local training:
$f_{n}(\theta_{g_{e}}^{n})=f_{n}(\theta_{g_{e}}^{n})+\frac{\mu}{2}||\theta_{g_{e}}^{n}-\theta_{g_{e}}||^{2},$ (1)
where $f_{n}(\cdot)$ is the loss function for edge device $n$, $\theta^{g_{e}}$ is the global model, and $\mu$ is a tunable parameter controlling the strength of the proximal term. After local training, the central server aggregates the updated models as in FedAvg. FedProx maintains stability across diverse client data distributions with the help of the proximal term.
FedRS: FedRS tackles data heterogeneity problem by scaling down updates for classes missing in an edge device’s local dataset. During local training, FedRS adjusts the predicted probability $\pi_{i,y}$ for class $y$ of sample $i$ by applying a scaling factor to the weights of missing classes as follows:
$\pi_{i,y}=\frac{\exp(\alpha_{y}\,w_{y}^{T}\,h_{i})}{\sum_{j=1}^{\mathcal{C}}\exp(\alpha_{j}\,w_{j}^{T}\,h_{i})},$ (2)
where $w_{y}$ is the weight vector associated with class $y$, and $h_{i}$ represents the feature vector for sample $i$. Here, $\alpha_{y}=1$ for observed classes and $\alpha_{y}=\alpha\in[0,1]$ for missing classes, effectively downscaling the influence of those classes. By limiting the effect of missing classes on the model update, this adjustment improves convergence in Non-IID settings when the central server aggregates the models as in FedAvg.
III Problem and Model
III-A Federated Heterogeneous Setting
In this section, we describe the mathematical objective of the FL setting and introduce our proposed plugin to handle heterogeneous data distributions across edge devices. We consider a federated network composed of $N$ edge devices, each communicating with a central server. Each edge device $n=1,\ldots,N$ holds a private dataset $\mathcal{D}^{n}$, with each dataset varying in both size and class distribution. Furthermore, edge devices possess different computational capabilities, affecting the number of local epochs $1\leq l_{e}\leq L$ within maximum value $L$, they can perform during each update round. FL aims to optimize a shared global model $\theta_{g_{e}}\in\Theta$ within the parameter space $\Theta$ without transferring the local datasets $\mathcal{D}=\cup_{n=1}^{N}\mathcal{D}^{n}$ to the central server, thus preserving data privacy.
To achieve this, the FL objective aims to minimize the global loss function $F(\theta)$, defined as:
$\min_{\theta_{g_{e}}\in\Theta}F(\theta)\triangleq\frac{1}{N}\sum_{n=1}^{N}f_{n}(\theta_{g_{e}}^{n}),$ (3)
where $F(\theta_{g_{e}})$ represents the global objective at the central server, and each local objective function $f_{n}(\theta)$ for edge device $n$ is defined as:
$f_{n}(\theta_{g_{e}}^{n})=\frac{1}{|\mathcal{D}^{n}|}\sum_{(x,y)\in\mathcal{D}^{n}}f_{n}(\theta_{g_{e}};x;y).$ (4)
Here, $f_{n}(\theta;x)$ represents the loss function for data sample $x$ within the local dataset $\mathcal{D}^{n}$. Due to the Non-IID nature, the distribution of data on edge devices $i$ and $j$ may differ significantly for $i\neq j$. Such statistical heterogeneity can lead to divergent local models, motivating the need for techniques to balance the data distribution, as proposed in this work.
III-B Federated learning with Proposed Plugin
In this section, we describe the integration of the proposed plugin with the standard FedAvg algorithm. Our proposed plugin consists of two phases: Phase 1, Data Augmentation with Generative AI, and Phase 2, Balanced Sampling as illustrated in Fig. 2. In the following, we present an example of applying FedAvg with the proposed plugin.
Figure 2: Illustration of the proposed plugin integrated into FL process, showing phase 1: data augmentation at edge devices and phase 2: balanced sampling at central server.
III-B1 Phase 1-Data Augmentation with Generative AI
As shown in Fig. 3, we utilize generative AI to alleviate computational and data imbalance on edge devices, which are representative challenges in FL. By generating synthetic data, especially for underrepresented classes, the technique augments the datasets, enabling a transformation from Non-IID distributions to approximate IID conditions. Providing a practical solution for edge devices with limited or skewed datasets, this data synthesis method helps to achieve a more balanced learning environment.
Figure 3: Illustration of data augmentation at edge device using generative AI to approximate Non-IID data to IID distributions. Original data is supplemented with synthetic data generated by generative AI to balance class distributions.
Let $\mathcal{D}^{n}=\{(x_{i}^{n},y_{i}^{n})\}_{i=1}^{|\mathcal{D}^{n}|}$ represent the local dataset at edge device $n$, where $x_{i}^{n}$ is the input data, and $y_{i}^{n}$ is the corresponding label. When certain classes are underrepresented in $\mathcal{D}^{n}$, a generative AI $\mathcal{G}$ is used to synthesize additional samples to balance the dataset:
$\mathcal{D}^{n}_{\text{aug}}=\mathcal{D}^{n}\cup\mathcal{G}(\mathcal{D}^{n}),$ (5)
where $\mathcal{G}(\mathcal{D}^{n})$ generates synthetic samples to balance the distribution. Specifically, the generative AI model produces samples for each class $y$ where data is insufficient, ensuring that class imbalance is minimized. To avoid introducing unseen data, the process only generates data for classes that are present in $\mathcal{D}^{n}$:
$\mathcal{G}(\mathcal{D}^{n})=\bigcup_{y\in\mathcal{C},|\mathcal{D}^{n}_{y}|>0}\mathcal{G}_{y}(\mathcal{D}^{n}),$ (6)
where $\mathcal{G}_{y}(\mathcal{D}^{n})$ denotes the synthetic data generated for class $y$, and $\mathcal{C}$ represents all classes.
To determine the quantity of synthetic data, we calculate the deficiency ratio for each class $y$ by comparing its size $|\mathcal{D}^{n}_{y}|$ with the maximum class size $|\mathcal{D}^{n}_{\text{max}}|$:
$\Delta_{y}=\frac{|\mathcal{D}^{n}_{\text{max}}|-|\mathcal{D}^{n}_{y}|}{|\mathcal{D}^{n}|},\quad\text{if }|\mathcal{D}^{n}_{y}|>0.$ (7)
If $\Delta_{y}$ is positive, indicating class imbalance, the generative model synthesizes additional samples as:
$\mathcal{G}_{y}(\mathcal{D}^{n})=\begin{cases}r(1,\Delta_{y}\cdot|\mathcal{D}^{n}|)&\text{if }|\mathcal{D}^{n}_{y}|>0\\ 0&\text{if }|\mathcal{D}^{n}_{y}|=0\end{cases}$ (8)
where $r(1,\Delta_{y}\cdot|\mathcal{D}^{n}|)$ is a random integer between 1 and $\Delta_{y}\cdot|\mathcal{D}^{n}|$. The final augmented dataset for each edge device $n$ is thus:
$\mathcal{D}^{n}_{\text{aug}}=\mathcal{D}^{n}\cup\bigcup_{y\in\mathcal{C}}\mathcal{G}_{y}(\mathcal{D}^{n}).$ (9)
Data augmentation ensures that the local model trained on $\mathcal{D}^{n}_{\text{aug}}$ benefits from a more balanced class distribution, improving the global model’s performance by mitigating the effects of Non-IID data. Consequently, this approach allows edge devices to contribute more effectively to the federated learning process, enhancing convergence speed and model robustness.
III-B2 Phase 2-Balanced Sampling
Following data augmentation phase, the central server implements a balanced sampling strategy as shown in Fig 4. This strategy optimizes model aggregation at the central server by focusing on edge devices whose set of data distributions most closely approximate an IID target. By prioritizing balanced data across edge devices, the strategy improves convergence, ensuring that model updates are derived from a representative subset of $\mathcal{D}$.
Figure 4: Illustration of the proposed balanced sampling at the central server, selecting set of $K=2$ edge devices with statistically representative data distributions for aggregation to achieve a optimized global model.
For each edge device $n$ with augmented dataset $\mathcal{D}^{n}_{\text{aug}}$, the proportion of data for each class $y$ is calculated as:
$p_{n}(y)=\frac{|\{x\in\mathcal{D}^{n}_{\text{aug}}:y=y\}|}{|\mathcal{D}^{n}_{\text{aug}}|},$ (10)
where $p_{n}(y)$ denotes the proportion of class $y$ in edge device $n$’s dataset. By collecting these class proportions from each device, the central server calculates a weighted global distribution $p_{\text{global}}(y)$ as:
$p_{\text{global}}(y)=\frac{\sum_{i}p_{i}(y)\cdot|\mathcal{D}^{i}_{\text{aug}}|}{\sum_{i}|\mathcal{D}^{i}_{\text{aug}}|}.$ (11)
Aiming for uniform class representation, the target IID distribution is:
$p_{\text{IID}}(y)=\frac{1}{|\mathcal{C}|},$ (12)
where $|\mathcal{C}|$ is the total number of classes. By measuring the distance between each edge device’s class distribution $p_{i}(y)$ and the target IID distribution $p_{\text{IID}}(y)$ using a distance function $d(\cdot)$, such as Euclidean or Manhattan distance, the central server assesses similarity to the target distribution:
$d(p_{i}(y),p_{\text{IID}}(y))=\|p_{i}(y)-p_{\text{IID}}(y)\|_{2}.$ (13)
Next, the central server then forms a subset by selecting K edge devices that show the smallest distances to the target IID distribution. Only these selected edge devices participate in the local learning stage, ensuring that the aggregation process draws from a well-balanced subset of data distributions. Each selected edge device performs local training on its augmented dataset $\mathcal{D}^{k}_{\text{aug}}$ using the AdamW [14] optimizer.
The local objective function for each selected edge device is defined as:
$f_{k}(\theta_{g_{e}}^{k})=\frac{1}{|\mathcal{D}^{k}_{\text{aug}}|}\sum_{(x_{i}^{k},y_{i}^{k})\in\mathcal{D}^{k}_{\text{aug}}}f_{k}(\theta_{g_{e}};x_{i}^{k},y_{i}^{k}),$ (14)
where $f_{k}(\theta_{g_{e}};x_{i}^{k},y_{i}^{k})$ represents the loss function (e.g., cross-entropy loss) for the sample $(x_{i}^{k},y_{i}^{k})$. By incorporating both adaptive moment estimation and weight decay, the AdamW optimizer updates the local model $\theta_{g_{e}}^{k}$, enhancing convergence and generalization.
AdamW updates the parameters as follows. First, compute the gradient of the local objective function:
$\textsl{g}_{l_{e}}=\nabla_{\theta_{g_{e}}}f_{k}(\theta_{g_{e},l_{e}}),$ (15)
where $l_{e}$ denotes the local epoch step. Update the first moment estimate $m_{l_{e}}$ and the second moment estimate $v_{l_{e}}$ using exponential moving averages:
$m_{l_{e}}$ $=\beta_{1}m_{l_{e}-1}+(1-\beta_{1})\textsl{g}_{l_{e}},$ (16)
$v_{l_{e}}$ $=\beta_{2}v_{l_{e}-1}+(1-\beta_{2})\textsl{g}_{l_{e}}^{2},$ (17)
where $\beta_{1}$ and $\beta_{2}$ are hyperparameters controlling the decay rates (e.g., $\beta_{1}=0.9$, $\beta_{2}=0.999$). Bias-corrected estimates are computed to adjust for the initialization bias:
$\hat{m}_{l_{e}}$ $=\frac{m_{l_{e}}}{1-\beta_{1}^{l_{e}}},$ (18)
$\hat{v}_{l_{e}}$ $=\frac{v_{l_{e}}}{1-\beta_{2}^{l_{e}}}.$ (19)
Update the model parameters with weight decay $\lambda$ as:
$\theta_{g_{e},l_{e}+1}=\theta_{g_{e},l_{e}}-\eta_{l}\left(\frac{\hat{m}_{l_{e}}}{\sqrt{\hat{v}_{l_{e}}}+\epsilon}+\lambda\theta_{g_{e},l_{e}}\right),$ (20)
where $\eta_{l}$ is the local learning rate, $\epsilon$ is a small constant to prevent division by zero, and $\lambda$ is the weight decay coefficient. Each edge device performs $1\leq l_{e}\leq L$ local epochs of training using these update rules, resulting in an updated local model $\theta^{k}_{g_{e}}$. By benefiting from the adaptive learning rates of AdamW and regularization through weight decay, the local optimization process leads to improved performance on heterogeneous data.
After local training, the central server aggregates the locally updated models $\vartheta=\{\theta^{k}_{g_{e}}\}_{k=1}^{K}$to update the next global model $\theta_{g_{e}+1}$. By aggregating the local models using weighted averaging based on dataset sizes, the global model is then updated:
$\theta_{g_{e}+1}=\frac{\eta_{g}}{K}\sum_{k=1}^{K}\theta_{g_{e}}^{k},$ (21)
where $\eta_{g}$ is the global learning rate. This process is repeated for each global epoch $g_{e}$, iteratively refining the global model until the desired performance is achieved or the maximum number of global epochs $G$ is reached. Overall procedure is described in Algorithm 1.
1
2Initialize: Global model $\theta_{0}$, maximum epochs $G$
3
4for global epoch $g_{e}=0$ to $G-1$ do
       /* Phase 1: Data Augmentation on Edge Devices */
5       for each edge device $n=1$ to $N$ in parallel do
6             for each class $y$ where $|\mathcal{D}^{n}_{y}|>0$ do
7                   Compute deficiency ratio $\Delta_{y}$
8                   if $\Delta_{y}>0$ then
9                         Generate synthetic data $\mathcal{G}_{y}(\mathcal{D}^{n})$
10
11
12            Form augmented dataset $\mathcal{D}^{n}_{\text{aug}}$
13
      /* Phase 2: Balanced Sampling at Central Server */
14       for each device $n=1$ to $N$ do
15             Compute class proportions $p_{n}(y)$
16             Compute distance $d_{n}=\|p_{n}(y)-p_{\text{IID}}(y)\|_{2}$
17
18      Select $K$ edge devices with smallest $d_{n}$
19       Broadcast: Send $\theta_{g_{e}}$ to selected $K$ devices
       /* Local Training on Selected Edge Devices */
20       for each device $s_{k}\in\mathcal{S}$ in parallel do
21             Initialize $\theta_{g_{e}}$ to local model
22             for local epoch $l_{e}=1$ to $L$ do
23                   Compute gradient $\textsl{g}_{l_{e}}=\nabla f_{k}(\theta^{k}_{g_{e},l_{e}})$
24                   Update $m_{l_{e}}$ and $v_{l_{e}}$ using $\textsl{g}_{l_{e}}$
25                   Compute $\hat{m}_{l_{e}}$ and $\hat{v}_{l_{e}}$
26                   Update model parameters
27                   $\theta_{g_{e},l_{e}+1}=\theta_{g_{e},l_{e}}-\eta_{l}\left(\frac{\hat{m}_{l_{e}}}{\sqrt{\hat{v}_{l_{e}}}+\epsilon}+\lambda\theta_{g_{e},l_{e}}\right)$
28
29            Obtain updated model $\theta^{k}_{g_{e}}$
30
      /* Aggregation at Central Server */
31       Receive models $\{\theta^{k}_{g_{e}}\}_{k=1}^{K}$
32       Update $\theta_{g_{e}+1}=\frac{\eta_{g}}{K}\sum_{k=1}^{K}\theta^{s_{k}}_{g_{e}}$
33
34Output: Final global model $\theta_{G}$
35
Algorithm 1 FL with Proposed Plugin
IV Experiment and Results
IV-A Experiment Setting
To validate the performance of our proposed balanced sampling and generative AI-based data augmentation approach, we conducted experiments on a medical text classification task [15]. Intel’s Gaudi 2 AI accelerator was used for its exceptional capability to handle the intensive computational demands of this process, ensuring efficiency and scalability in our experiments. The classification models were trained using a Non-IID data distribution across $N=100$ edge devices. In our experiments, we set the number of selected edge devices participating in each global round to $K=10$. We applied the proposed method to enhance the performance of commonly used federated learning algorithms, including FedProx and FedRS, which are designed to handle heterogeneous environments in FL.
For data augmentation, we utilized the Ollama API to leverage representative Large Language Models (LLMs) to generate synthetic data from original data. We used multiple LLMs for data generation, including Gemma 1.0 (7B parameters) [16], Gemma 2.0 (2B and 9B parameters) [17], Llama 3.1 (8B), Llama 3.2 (1B and 3B) [18], Phi 2 (3B) [19], Phi 3 (3B), and Phi 3.5 (3B) [20]. To apply the same synthetic data across various experiments, we pre-generated the data using these LLMs and loaded them into our code. By employing these models, we augmented the underrepresented classes in the datasets of the edge devices, thereby balancing the class distributions and mitigating the effects of Non-IID data.
To introduce statistical heterogeneity, we used a Dirichlet distribution to partition the overall training dataset among the edge devices, simulating Non-IID conditions across the devices [21]. Each device was assigned a portion of the dataset according to a Dirichlet distribution with concentration parameter $\alpha$, where lower values of $\alpha$ increase the degree of label skew. Computational heterogeneity was modeled by randomly selecting the number of local update epochs $l_{e}$ for each device, with $l_{e}$ values ranging between 1 and 5 for every global epoch $g_{e}$. Further details on hyperparameter settings and experimental configurations are available in the open-source repository [^1] https://github.com/NAVER-INTEL-Co-Lab/gaudi-fl.
IV-B Results
IV-B1 Impact on AI Model Diversity
Figure 5: Performance comparison of text classification models with and without our proposed plugin. Results demonstrate significant accuracy improvements across all models, with particularly notable gains in domain-specialized models (BioBERT, BioMedBERT) compared to their baseline performance.
In this experiment, we examined how our proposed plugin enhances the performance of various text classification models in FL environments. We compared performance improvements for four different models: general-purpose models BERT [22] and DistilBERT [23], and medical domain pre-trained models BioBERT [24] and BioMedBERT [25]. Through this experimental setup, we aimed to demonstrate our plugin’s capability to enhance model performance regardless of their initial domain specialization. Additionally, for data augmentation in Phase 1, multiple LLMs were utilized to generate synthetic data, ensuring balanced representation across all edge devices.
Performance evaluation revealed significant improvements across all models when applying our proposed plugin. As shown in Fig. 5, BERT’s accuracy increased from 29.8% to 47.3% and DistilBERT improved from 32.2% to 46.6%. Most notably, medical domain pre-trained models showed even more substantial gains, with BioBERT’s accuracy rising from 42.1% to 53.4% and BioMedBERT reaching 52.0% from a baseline of 28.7%. These significant improvements demonstrate our plugin’s effectiveness in enhancing model performance while preserving the advantages of domain-specific pre-training. The results particularly highlight how our plugin can amplify the benefits of domain-specialized models in federated medical applications. Based on these results, we use BioBERT as the edge device AI model for subsequent experiments.
Detailed analysis revealed the synergistic effect between our plugin and domain-specialized models. The performance gains were particularly pronounced for medical domain pre-trained models, with BioBERT showing 7.7% and 6.8% higher accuracy compared to BERT and DistilBERT respectively when enhanced by our plugin. These improvements demonstrate that our synthetic data augmentation and balanced sampling strategies effectively complement domain-specific knowledge, creating more robust FL systems. Furthermore, the significant performance enhancements across all models validate our plugin’s versatility in addressing data heterogeneity challenges.
IV-B2 Impact on Various FL Methods
Figure 6: Performance comparison of different FL algorithms enhanced with the proposed plugin. Results demonstrate consistent improvements in test accuracy across FedAvg, FedProx, and FedRS, with various LLM configurations, highlighting the plugin’s effectiveness and versatility.
To demonstrate the versatility and effectiveness of our proposed plugin across different federated optimization techniques, we designed comparative experiments. Our evaluation framework considered three key aspects: the plugin’s adaptability to various FL algorithms, its impact on convergence speed, and overall accuracy improvements. Through this comprehensive evaluation setup, we aimed to validate that our plugin can enhance performance across different FL approaches.
Performance analysis revealed consistent improvements across all FL algorithms when combined with our proposed plugin. Among tested combinations, FedRS with Llama3.0:8B achieved the highest accuracy at 53.67%, while the same configuration with FedProx and FedAvg reached 53.51% and 53.42% respectively as shown in Fig. 6. Notably, all three FL algorithms showed significant performance gains compared to their baseline implementations, with improvements ranging from 9% to 11%. This consistent enhancement across different FL approaches demonstrates the plugin’s effectiveness in mitigating data heterogeneity through synthetic data augmentation and balanced sampling.
Convergence analysis showed that all configurations with our plugin reached stable performance significantly faster than baseline approaches. While baseline FL algorithms required 23 epochs to reach maximum accuracy, configurations with our plugin consistently achieved convergence between epochs 15-20. Specifically, Gemma 2.0:9B, Llama 3.2 series, and Phi 2.0:3B demonstrated rapid convergence at epoch 15 across all FL algorithms, showing approximately 35% reduction in required training epochs. These results suggest that our proposed plugin effectively accelerates convergence regardless of the underlying FL algorithm or LLM choice. The experimental outcomes demonstrate that our synthetic data augmentation and balanced sampling strategies not only improve accuracy but also significantly enhance training efficiency across different FL frameworks.
IV-B3 Impact of Data Heterogeneity
Figure 7: Performance comparison across different degrees of data heterogeneity $\alpha$. Results demonstrate the plugin’s effectiveness in maintaining high accuracy even under extreme Non-IID conditions, with consistent improvements observed across various LLM configurations.
To validate our plugin’s effectiveness under different levels of data heterogeneity, we conducted experiments with varying degrees of Non-IID distribution controlled by the concentration parameter $\alpha$. Lower values of $\alpha$ indicate higher data heterogeneity among edge devices, presenting more challenging FL scenarios. This experimental setup particularly focuses on extreme Non-IID situations where conventional federated learning approaches often struggle to maintain performance.
Performance analysis revealed that our proposed plugin demonstrates increasingly significant improvements as data heterogeneity increases. With Gemma 1.0:7B, the plugin maintained 38.89% accuracy even at $\alpha=0.01$, while baseline FedAvg accuracy dropped to 21.43% as shown in Fig. 7. Similar patterns were observed across different LLM families, with Gemma 2.0:2b and Gemma 2.0:9b achieving 48.13% and 47.68% accuracy respectively at high heterogeneity levels. These results indicate that synthetic data generation effectively compensates for data imbalances, particularly in scenarios with minimal data overlap between devices.
Figure 8: UMAP visualization of embedding vectors from original and synthetic data across different LLM configurations. Distinct clustering patterns reveal varying capabilities of LLMs in preserving the original data distribution, with Phi 3.0 and 3.5 showing notable deviations from the source distribution.
Notably, under extreme Non-IID conditions $\alpha\leq 0.001$, our plugin consistently provided substantial performance improvements compared to unused case. The enhanced performance remained stable across all tested LLM configurations, with accuracy improvements ranging from 11% to 28% at $\alpha=0.001$. These findings demonstrate that our plugin effectively addresses one of the fundamental challenges in FL: maintaining model performance under severe data heterogeneity. The robust performance across different heterogeneity levels suggests strong potential for real-world applications where Non-IID data distribution is inevitable.
IV-B4 Analysis of Synthetic Data Distribution
To understand the effectiveness of different LLMs in generating representative synthetic data, we conducted an embedding space analysis. We visualized the distribution of both original and synthetic samples using UMAP [26] dimensionality reduction technique as shown in Fig. 8. Embedding vectors were extracted using ’MXBai-Embed-Large model’, providing a two-dimensional representation of the high-dimensional text features. This visualization approach enables direct comparison of the semantic relationships between original medical texts and their synthetic counterparts. Such comparisons offer insights into each LLM’s capability to maintain the underlying data distribution.
Comparative analysis revealed distinct patterns across different LLM families in preserving the original data distribution. Among them, Gemma 2.0:9B and Llama 3.2:3B demonstrated particularly strong alignment between synthetic and original data clusters, maintaining clear separation between different medical text categories while preserving intra-class relationships. By successfully capturing the inherent structure of the medical text data, these models validated their superior performance observed in previous experiments. Additionally, the synthetic samples generated by these models showed consistent overlap with original data clusters, providing clear evidence for their effective preservation of domain-specific features.
Notably, Phi 3.0:3B and Phi 3.5:3B exhibited significant deviation from the original data distribution. Their synthetic samples formed distinct clusters that were clearly separated from the original data points, unlike other LLM families. This distribution mismatch explains their relatively lower performance in previous experiments, particularly in maintaining classification accuracy. This misalignment between original and synthetic data distributions particularly impacts the effectiveness of our plugin’s data augmentation strategy. These findings ultimately highlight the importance of selecting appropriate generative models that can maintain semantic consistency for synthetic data generation.
V Conclusion
In this paper, we proposed a novel plugin for FL that addresses data heterogeneity challenges through generative AI-enhanced data augmentation and balanced sampling strategy. Our experiments demonstrated the plugin’s effectiveness across various text classification models and FL algorithms while reducing required training epochs. The plugin showed strong performance improvements in medical domain-specialized models, with BioBERT achieving 53.4% accuracy from a 42.1% baseline. Our approach maintained robust performance under extreme Non-IID conditions, showing consistent improvements across different LLM configurations. These results establish our plugin as a flexible solution for enhancing FL performance in real-world heterogeneous environments.